EVA Report – December 29th

Request: Crew 185 EVA #9 for SOL 11.
Propose of EVA:
-David did not sleep well, so he did not take part to the EVA. Only Ilaria and John were on EVA.
-We collected only one sample at Yellow Moon.
-We came back just after.

EVA #9:
Start time: 1030am
End time: 1330pm -> 12:20 pm

Ilaria: Leader
John: Astros
Arno, David and Thibo are Comms.

Destination: Yellow Moon (38.4 N – 110.82 W).
Vehicles: ATV 3 (John), ATV 3 (Ilaria)
Prepared by Ilaria Cinelli

Greenhab Report – December 29th

Greenhab Report
David Murray
29 Dec, 2017

Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with window/door open from 9:05 – 16:57
Working Hour: 14:21
Inside temp at working hour: 22 C
Outside temp during working hours: 5 C
Inside temperature high: 24 C
Inside temperature low: 16 C
Inside humidity: 23%

Heating
Functioning nominally
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A

Changes to Crops
None
Morning Research Observations
N/A
Daily Water Usage for Crops
6 gallons
Time(s) of Watering Crops
9:15
Aquaponics
N/A

Narrative
The GreenHab is clean and ready for the next crew. They’ll be inheriting a fertilizer sprayer, about seven hundred grams of moringa leaf and shoot matter, and a blender. All of the plants are healthy and no plants were lost unintentionally. I have left a note indicating when the next application of each fertilizer is due. Everything has been cleaned. Tomorrow evening I’ll have an opportunity to train the incoming GreenHab Officer and to meet the rest of Crew 186.

Support/Supplies Needed
Sand to be used as a substrate for moringa.

Sol Summary – December 29th

Crew 185 Sol Summary Report 29 Dec 2017

Sol 11

Last Day of Sim

David Murray

Mission Status: Everything is going well

Sol Activity Summary: We woke up at 0800, ate breakfast and got ready for the EVA. At 1034, the EVA crew entered the airlock and returned at 1218. While the EVA crew was gone, Arno, Tibo, and I cleaned the hab and the greenhab. Upon return we ate lunch, packed up our belongings, and continued cleaning. Later we took inventory of the remaining food and prepared MDRS for the next crew.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: None

Weather: Sunny, warm

Crew Physical Status: Healthy

EVA: Exploration, sample collection

Reports to be filed: Journalist Report, Greenhab Report, Sol Summary, EVA report

Support Requested: None

End of Mission Summary – Crew 185

Mars Desert Research Station

End of Mission Summary

Crew 185 – Mars Society Crew #2

International Crew

 

Crew 185 (Dec. 16th, 2017 – Dec 31st, 2017):

Commander: Ilaria Cinelli (Italy/Ireland)
Executive Officer/Crew Engineer: Thibault Paris (France)
Greenhab Officer/Crew Biologist: David Murray (United States)
Crew Engineer: Arno Passaron (France)
Crew Health & Safety Officer: John Sczepaniak (United States)

Today is the last day of SIM, and tomorrow morning my Crew and I will land on Earth!!!

As you may know I am Commander of an international Crew, selected by The Mars Society, and this is an international mission getting close to the end. I will mention a few of our activates to show you how much (international) science we can get out of two-week mission.

Yesterday, we have completed a very interesting experiment about “Shared Spatial Representation” of the environment around Astros on EVA and HabCom (in the Hab), during which we were looking for objects placed in specific places, guided by with the vocal indication of HabCom (so without GPS and tech), (PI IMS laboratory and Association Planète Mars, France).

Then, today we have also completed two undergraduate student projects about sediment movement by aeolian transport and response of the use of a loaded vehicle over a range of terrain types (PI Trinity College Dublin, Republic of Ireland).

We also did different surveys to evaluate the quality of life in the Hab. One of this was the use of a software to estimate the stress level during writing (PI Mars Planet – Italian Mars Society, Italy). Then, other are about crew dynamics. Pre-flight tests evaluate individual differences before the mission (PI University of Padua, Italy), and other tests evaluate individual performance during the mission (PI 100 Year Starship, USA). Instead, team cohesion was archived by practicing empowerment (PI Fondazione Internazionale verso l’Etica – FIVE – onlus, Italy) and mindfulness (PI University of London, UK), and table games!

Then, two studies about safety were pretty useful to develop awareness in isolation analysing both the context (PI William Carey College of Osteopathic Medicine, USA) and the Crew (PI Mars Planet – Italian Mars Society, Italy).

In the end, we did some fun activities about education and outreach. One was about the “Kid2Mars” project where children from all over the World asked questions about Mars (PI InnovaSpace). The other was about “Crea(c)tivity”, a two-day workshop during which secondary school students have been designing and engineering prototypes that can have real space applications (PI ISIA Firenze, Italy).

A very original experiment was about clothing and textiles understanding the needs and constraints of design for apparel and habitats for long-duration space exploration and habitation (PI University of Rhode Island and 100 Year Starship, USA).

Then, we are honoured to have tested the first prototype of a 3D printed spacesuit, called X-1 (PI Ecole polytechnique, France). This project was supported by the French chapter of the Mars Society (PI The Association Planete Mars) to develop and test this prototype.

Additionally, we have been utilizing the full potential of crops to imitate conditions found in a limited resource environment. By producing a fertilizer from the most nutritious plant on Earth, we not only get nutritional value from this plant but hopefully the ability to increase the yields of other crops.

In the end, crew 185 completed an important anesthesia task during the mission looking at the ability of astronauts to complete a nerve block of the lower leg. They used gel models created at the University of California, San Diego to place a needle above and below the simulated nerve located behind a knee.  The simulation looks at the ability of participants’ time to complete tasks in an emergency scenario.

I can say we had fun while working! There are so many things happening everyday that there is not worries to be bothered during the day (and night) at the Hab.

If it is not a technical problem, it will be a human factor issue within the Crew! I LOVE what I am doing, I will never get tired of these challenges! I really LOVE what I am doing, hard to tell you how much!

Really, thank you for having me as Commander, and thank you for this great Crew!

I am learning so much in management, communication and science, and I am trying to give back good quality data that can be use for on-going and future research projects. I am feeling so enriched that I need to share my empowerment! So, science is the best way though which I can have a positive impact in research.

Thus, thank you so much! This is a great personal experience, and I see my Crew getting the most out of it. It is time for me to let them go, I have trained them for the best and I believe these two weeks gave them enough experience to make them stronger in the future.

Again, thank you to all the Team of Mission Support and The Mars Society!

In particularly, thank you Shannon!

Commander Ilaria Cinelli

 

Special thanks to:

  • The Mars Society
  • Mission Support
  • The Mars Desert Research Station
  • IMS laboratory and Association Planète Mars, France.
  • Trinity College Dublin, Republic of Ireland.
  • Mars Planet – Italian Mars Society, Italy.
  • University of Padua, Italy.
  • 100 Year Starship, USA.
  • Fondazione Internazionale verso l’Etica – FIVE – onlus, Italy.
  • University of London, UK.
  • ISIA Firenze, Italy.
  • University of Rhode Island, USA.
  • The Association Planete Mars, France.
  • Ecole polytechnique, France.
  • William Carey College of Osteopathic Medicine, USA.
  • InnovaSpace.

The Mars Desert Research Station – Crew 185

 

Crew 185 – X-1 (PI Ecole polytechnique, France)

 

Crew 185 – Green Hab – Beans

Journalist Report – December 28th

Today, we had a lie-in and woke up at 8am. After the breakfast, we made sure to charge the generator, we played poker and prepared for the EVA. Ilaria, John and I then took soil samples on our ATVs and our rover. We hiked to the top of a hill and enjoyed the landscape. We were surrounded by a large canyon and desert plains. As we went back to the Hab, we did the last communication experiment of our mission. After this 3-hour-long EVA, we had a well-deserved lunch. Everybody went back to their daily activities: I cleaned the helmets and checked the electronic systems, David worked in the GreenLab, while Thibault, Ilaria and John did a medical experiment.

Arno

EVA Report – December 28th

Report: Crew 185 EVA #8 for SOL 10.

Propose of EVA:
-Collecting geological samples for soil density study. We will do multiple stops along Cow Dung road from the Hab to Gateway Lith. -We have done two stops along the way to collect samples, and it was a good teamwork!
-We stayed at the Lith for about 30 min for exploration and observing the panorama!
-Astro-CapCom communication. CapCom will drive the Astros to a specific geographical position (nearby the Hab) without the use of a GPS. EVA #8:
Start time: 1030am
End time: 1330pm->1430pm
Ilaria: Leader
Arno and John: Astros
David and Thibo are Comms.
Destination: Gateway to Lith (110.81 W – 38.45 N), see attachment. Vehicles: ATV (Arno), ATV (John), Rover Spirit (Ilaria)

Prepared by Ilaria Cinelli

GreenHab – December 28th

David Murray

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient
Ambient with window/door open from 9:40 – 18:06
Working Hour: 15:38
Inside temp at working hour: 23 C
Outside temp during working hours: 8 C
Inside temperature high: 26 C
Inside temperature low: 15 C
Inside humidity: 60%

Heating
Functioning nominally
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
N/A
Changes to Crops
None
Morning Research Observations
N/A
Daily Water Usage for Crops
6 gallons
Time(s) of Watering Crops
15:25

Aquaponics
N/A
Narrative
I replaced some of the tomato cages around the beans with wooden poles. Fourteen moringa seeds were planted in soil and five were started in damp paper towels.
Support/Supplies Needed
N/A

Sol Summary – December 28th

Sol 10

Long EVA

David Murray

Mission Status: Everything is going well

Sol Activity Summary: We woke up at 0800, had breakfast, and switched off the generator. Today, the EVA crew got a chance to gain experience on the ATV’s and to explore Lith Canyon. Then, they completed the navigation exercise and returned to the hab. After lunch, I planted Moringa oleifera seeds in the greenhab and watered the plants, Tibo and Ilaria participated in John’s nerve block experience, and Arno checked our state of charge. The crew plans to eat dinner and finish the rest of the movie that we started last night before going to bed.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: None

Weather: Sunny, warm

Crew Physical Status: Healthy

EVA: Exploration, sample collection

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary, EVA report

Support Requested: None

Sol Summary – December 27th

Sol 9

Ninth Day on Mars

David Murray

Mission Status: Everything is going well

Sol Activity Summary: The crew checked the generator, ate breakfast, and prepared for EVA. The EVA crew returned to the airlock to address complications with the radios. Once that was taken care of we completed the next half of the exercise. While we were finishing our EVA, we noticed a scratch in the helmet. We finished our task before returning to the airlock. After EVA, we completed our list of group tasks, wrote reports, and had dinner.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: Scratch noticed on helmet during EVA

Weather: Sunny

Crew Physical Status: Healthy

EVA: Navigation

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary, EVA report

Support Requested: None

Greenhab Report – December 27th

Greenhab Report

David Murray

27 Dec, 2017

Environmental control:

Ambient

Ambient with window/door open from 14:05 – 16:20

Working Hour: 14:14

Inside temp at working hour: 35 C

Outside temp during working hours: 7 C

Inside temperature high: 40 C

Inside temperature low: 19 C

Inside humidity: 46%

Heating

Functioning nominally

Turns on at 16 C and off at 20 C

Cooling

Not functional

Shade Cloth On/Off

On

The greenhab lights are on from 17:00 to 24:00 every night

Average temperature:

N/A

Changes to Crops

Cucumbers are wilting because of the heat.

Morning Research Observations

N/A

Daily Water Usage for Crops

6 gallons

Time(s) of Watering Crops

14:14

Aquaponics

N/A

Narrative

The cucumbers were wilting because the inside temperature got too hot and the soil was dry. They recieved extra water today and they should be fine by tomorrow.

Support/Supplies Needed

N/A